# Consistent estimation of generative model representations in the data kernel perspective space

### v1

@@ Abstract
Generative models, such as large language models and text-to-image diffusion models, produce relevant information when presented a query. Different models may produce different information when presented the same query. As the landscape of generative models evolves, it is important to develop techniques to study and analyze differences in model behaviour. In this paper we present novel theoretical results for embedding-based representations of generative models in the context of a set of queries. We establish sufficient conditions for the consistent estimation of the model embeddings in situations where the query set and the number of models grow.
@@ 1 Introduction
In this paper we provide theoretical justification for these successful empirical investigations by arguing that the induced perspective space is consistent for a population-level perspective space. In particular, we analyze three different – progressively more complicated – settings in which the perspective space can be used: i) fixed collection of models and fixed set of queries; ii) fixed collection of models and growing set of queries; and iii) growing collection of models and growing set of queries. For each setting we show that the multi-dimensional scaling with the raw stress criterion of a collection of matrix representations yields vectors that are consistent for an appropriately defined limiting configuration. Further, we provide sufficient conditions for consistency as a function of the generative properties of each model.
@@ 2.2 Multidimensional scaling by raw stress embedding
Multidimensional scaling ([3]) refers to the family of methods that produce vector representations of a set of objects from their pairwise dissimilarities. Raw stress embedding ([14, 22]) is a popular method for multidimensional scaling. For our purposes, we discuss raw stress embedding in the context of two scenarios: finite sample size and infinite sample size. The following is adapted from [21] for our specific setting of collections of generative models.
@@
Given dissimilarity matrix $\boldsymbol{\Delta}^{(\infty)}$, $\boldsymbol{\psi}=\mathrm{mds}(\boldsymbol{\Delta}^{(\infty)})$ is not necessarily unique. We denote the set of all solutions $\mathrm{mds}(\boldsymbol{\Delta}^{(\infty)})$ by $\mathrm{MDS}(\boldsymbol{\Delta}^{(\infty)})$.
@@
Note that $\min_{\mathbf{z}_{i}\in\mathbb{R}^{d}}\sum_{i,i^{\prime}=1}^{n}\left(\left\lVert\mathbf{z}_{i}-\mathbf{z}_{i^{\prime}}\right\rVert-\boldsymbol{\Delta}_{ii^{\prime}}^{(\infty)}\right)^{2}=\min_{\tilde{\boldsymbol{\Delta}}\in\mathcal{Y}_{n}}\left\lVert\tilde{\boldsymbol{\Delta}}-\boldsymbol{\Delta}^{(\infty)}\right\rVert_{F}^{2}$, where $\mathcal{Y}_{n}$ is the set of all $n\times n$ Euclidean distance matrices with Frobenius norm less than or equal to $\left\lVert\boldsymbol{\Delta}^{(\infty)}\right\rVert_{F}$, which is closed, bounded, and complete [22]. This guarantees the existence of a solution to Eq. (1).
@@
$\sigma((\boldsymbol{\Delta}^{(\infty)},\mathcal{P}),h)=\int_{\mathcal{M}}\int_{\mathcal{M}}\left(\left\lVert h(\mathbf{m}^{\prime})-h(\mathbf{m}^{\prime\prime})\right\rVert-\boldsymbol{\Delta}^{(\infty)}(\mathbf{m}^{\prime},\mathbf{m}^{\prime\prime})\right)^{2}\mathcal{P}(d\mathbf{m}^{\prime})\mathcal{P}(d\mathbf{m}^{\prime\prime}).$ (2)
@@
Note that the function $\tilde{\boldsymbol{\Delta}}:\mathcal{M}\times\mathcal{M}\to\mathbb{R}$ defined as $\tilde{\boldsymbol{\Delta}}(\mathbf{m^{\prime}},\mathbf{m^{\prime\prime}})=\left\lVert h(\mathbf{m^{\prime}})-h(\mathbf{m^{\prime\prime}})\right\rVert$ is an Euclidean pseudometric on $\mathcal{M}$ to $\mathbb{R}^{d}$. Since the set $\mathcal{Y}$ of all Euclidean pseudometrics from $\mathcal{M}$ to $\mathbb{R}^{d}$ is closed and complete, $\sigma((\boldsymbol{\Delta}^{(\infty)},\mathcal{P}),h)$ can be minimized by varying $h$ over the set of all Borel-measurable embedding functions from $\mathcal{M}\to\mathbb{R}^{d}$.
@@ 3 Description of the setting
We consider a generative model $f$ to be a (random) mapping from a query space $\mathcal{Q}$ to an output space $\mathcal{X}$. Let $f_{1},\dots f_{n}$ be $n$ generative models with shared query and output spaces and $\{q_{1},\dots q_{m}\}$ be a collection of $m$ queries. For our purposes, every model $f_{i}$ responds to every query $q_{j}$ exactly $r$ times and we let $f_{i}(q_{j})_{k}$ denote the $k$-th replicate of the response of $f_{i}$ to $q_{j}$. We assume that there exists an embedding function $g$ that maps a response to a vector in $\mathbb{R}^{s}$. For every (model, query) pair, we let $F_{ij}$ be the distribution of embedded responses and assume $g(f_{i}(q_{j})_{k})\sim^{iid}F_{ij}$.
Let $\bar{\mathbf{X}}_{i}$ be the $m\times s$ matrix whose $j^{th}$ row is the mean over replicates of the $i^{th}$ model’s response to the $j^{th}$ query: $(\bar{\mathbf{X}}_{i})_{j\cdot}=\frac{1}{r}\sum_{k=1}^{r}g(f_{i}(q_{j})_{k})\approx(\boldsymbol{\mu}_{i})_{j\cdot}\equiv\mathbb{E}_{F_{ij}}\left[g(f_{i}(q_{j})\right]$.
Further, let $\mathbf{D}$ be the $n\times n$ pairwise distance matrix with entries $\mathbf{D}_{ii^{\prime}}=\frac{1}{m}||\bar{\mathbf{X}}_{i}-\bar{\mathbf{X}}_{i^{\prime}}||_{F}$. Then, as $r\to\infty$, $D_{ii^{\prime}}$ converges to the model mean discrepancy $\boldsymbol{\Delta}_{ii^{\prime}}=\frac{1}{m}||\boldsymbol{\mu}_{i}-\boldsymbol{\mu}_{i^{\prime}}||_{F}$.
@@ 4.2 Fixed set of models and growing set of queries
For some $q\in\mathbb{N}$, there exist vectors $\boldsymbol{\phi}_{1},\dots\boldsymbol{\phi}_{n}\in\mathbb{R}^{q}$, such that for every pair $(i,i^{\prime})\in[n]^{2}$, $\boldsymbol{\Delta}_{ii^{\prime}}=\frac{1}{m}||\boldsymbol{\mu}_{i}-\boldsymbol{\mu}_{i^{\prime}}||\to\left\lVert\boldsymbol{\phi}_{i}-\boldsymbol{\phi}_{i^{\prime}}\right\rVert$ as $r\to\infty$.
@@ 4.3 Growing set of models and growing set of queries
Figure 1: Numerical evidence of the consistency of $\widehat{\boldsymbol{\psi}}$ to $\boldsymbol{\psi}$ for fixed $n=10$ and fixed $m=50$. The left figure shows the 2-d DKPS of $10$ models induced with $20$ queries, $50$ replicates each, and the embedding function NomicEmbedv1.5. The red circles have radius equal to the average $L_{2}$ between $\widehat{\psi}^{*}$ and model representations estimated with $r\in\{1,5,20\}$ replicates for each query. The right figure shows the distribution of the average $L_{2}$ norm in the DKPS across models for various $r$.
@@ 5 Numerical experiments
We next provide empirical support of the consistency of the representations of models in the three settings we analyzed above. For each setting, we use LLaMA-3-8B-Chat [8] as a base model and consider up to $50$ models. The models are parameterized by different fixed context augmentations $a_{i}$, i.e., $f_{i}=f(\;\cdot\;;a_{i})$. Each $a_{i}$ is a text string related to RA Fisher such as $a_{i}=$ “RA Fisher pioneered the principles of the design of experiments.” or $a_{i^{\prime}}=$ “RA Fisher’s view on eugenics were primarily based on anecdotes and prejudice." written by us. Given a query $q_{j}$, the base model is prompted with the appropriately formatted prompt ``$a_{i}\;q_{j}"$ $50$ times. In our experiments we consider up to $200$ questions about RA Fisher such as $q_{j}=$ “What is R.A. Fisher’s most well-known statistical theorem?" as queries. The questions were sampled from ChatGPT with the prompt “Provide 200 questions related to RA Fisher". We use NomicEmbedv1.5 [17] as the embedding model and GrasPy’s [6] implementation of multi-dimensional scaling throughout.
@@
We do not know the “ground truth" DKPS $\boldsymbol{\psi}$ in any setting, and thus treat $\widehat{\boldsymbol{\psi}}$ estimated with $r=50$ as a proxy for the ground truth. For settings with growing $n$ and/or growing $m$, we similarly treat $\widehat{\boldsymbol{\psi}}^{(50)}$ estimated with $n=50$ and $m=200$ as the ground truth. We refer to these estimates as $\widehat{\boldsymbol{\psi}}^{*}$. We let $d=2$ in all settings for visualization purposes. Finally, in settings where $n,m,$ or $r$ grow, we sample from the appropriate set of models, queries, or replicates with replacement to calculate $\widehat{\boldsymbol{\psi}}$ and report the average $L_{2}$ norm or average two-to-infinity norm of the difference between $\boldsymbol{\widehat{\psi}}^{*}$ and $\boldsymbol{\widehat{\psi}}\mathbf{W}$ of 10 bootstrap samples, where $\mathbf{W}\in\mathcal{O}(2)$ is the Procrustes solution [11].
@@ 5.1 Fixed collection of models and fixed set of queries
As with the theoretical analysis, our empirical analysis starts in the simplest setting where $n=10$ and $m=20$ are fixed while $r$ grows. The collection of $10$ models and set of $20$ queries were selected at random without replacement. The target DKPS $\boldsymbol{\widehat{\psi}}^{*}$ is shown on the left of Figure 1. Each black dot represents a model. The red circles have radius equal to the average (across bootstrap samples) $L_{2}$ norm between $\boldsymbol{\widehat{\psi}}^{*}$ and $\boldsymbol{\widehat{\psi}}^{(r)}\mathbf{W}$ for selected $r$. The right figure of Figure 1 shows the distribution of the average (across bootstrap samples) $L_{2}$ norm for each model for more values of $r$. The decreasing average $L_{2}$ norm for all models as $r$ increases supports Theorem 1.

### v2

@@ Abstract
Generative models, such as large language models and text-to-image diffusion models, produce relevant information when presented a query. Different models may produce different information when presented the same query. As the landscape of generative models evolves, it is important to develop techniques to study and analyze differences in model behaviour. In this paper we present novel theoretical results for embedding-based representations of generative models in the context of a set of queries. In particular, we establish sufficient conditions for the consistent estimation of the model embeddings in situations where the query set and the number of models grow.
@@ 1 Introduction
In this paper we provide theoretical justification for these successful empirical investigations by arguing that the estimated perspective space is consistent for a population-level perspective space. In particular, we analyze three different – progressively more complicated – settings in which the perspective space can be used: i) fixed collection of models and fixed set of queries; ii) fixed collection of models and growing set of queries; and iii) growing collection of models and growing set of queries. For each setting we show that the multi-dimensional scaling with the raw stress criterion of a collection of matrix representations yields vectors that are consistent for an appropriately defined limiting configuration. Importantly, we provide sufficient conditions for consistency as a function of the generative properties of each model.
@@
Contribution. We build upon recent theoretical results related to multi-dimensional scaling via the raw stress criterion. In particular, we provide sufficient conditions for the consistency of sampling-based vector representations of black-box generative models. Our results are general to collections of generative models whose outputs can be mapped to a shared embedding space.
@@ 2.2 Multidimensional scaling by raw stress embedding
Multidimensional scaling [3] refers to the family of methods that produce vector representations of a set of objects from their pairwise dissimilarities. Raw stress embedding [15, 23] is a popular method for multidimensional scaling. For our purposes, we discuss raw stress embedding in the context of two scenarios: finite sample size and infinite sample size. The following is adapted from [23] for our specific setting of collections of generative models.
@@
Given dissimilarity matrix $\boldsymbol{\Delta}^{(\infty)}$, $\boldsymbol{\psi}=\mathrm{mds}(\boldsymbol{\Delta}^{(\infty)})$ is not unique – an affine transformation upon a minimizer gives another minimizer. We denote the set of all solutions $\mathrm{mds}(\boldsymbol{\Delta}^{(\infty)})$ by $\mathrm{MDS}(\boldsymbol{\Delta}^{(\infty)})$. Apart from affine transformations of a solution, there may be two different minimizers that are not affine transformation of each other.
@@
Note that
$\min_{\mathbf{z}_{i}\in\mathbb{R}^{d}}\sum_{i,i^{\prime}=1}^{n}\left(\left\lVert\mathbf{z}_{i}-\mathbf{z}_{i^{\prime}}\right\rVert-\boldsymbol{\Delta}_{ii^{\prime}}^{(\infty)}\right)^{2}=\min_{\tilde{\boldsymbol{\Delta}}}\left\lVert\tilde{\boldsymbol{\Delta}}-\boldsymbol{\Delta}^{(\infty)}\right\rVert_{F}^{2},$
where $\tilde{\boldsymbol{\Delta}}$ is varied over the set of all $n\times n$ Euclidean distance matrices. From the proof of Theorem 2 in [23], it can be argued that it is enough to vary $\tilde{\boldsymbol{\Delta}}$ over $\mathcal{Y}_{n}$, the set of all $n\times n$ Euclidean distance matrices with Frobenius norm less than or equal to $\left\lVert\boldsymbol{\Delta}^{(\infty)}\right\rVert_{F}$, which is closed, bounded, and complete. This guarantees the existence of a solution to Eq. (1).
@@
$\sigma((\boldsymbol{\Delta}^{(\infty)},\mathcal{P}),h)=\int_{\mathcal{M}}\int_{\mathcal{M}}\left(\left\lVert h(\mathbf{m}^{\prime})-h(\mathbf{m}^{\prime\prime})\right\rVert-\boldsymbol{\Delta}^{(\infty)}(\mathbf{m}^{\prime},\mathbf{m}^{\prime\prime})\right)^{2}\mathcal{P}(d\mathbf{m}^{\prime})\mathcal{P}(d\mathbf{m}^{\prime\prime}).$
@@
The function $\tilde{\boldsymbol{\Delta}}:\mathcal{M}\times\mathcal{M}\to\mathbb{R}$ defined as $\tilde{\boldsymbol{\Delta}}(\mathbf{m^{\prime}},\mathbf{m^{\prime\prime}})=\left\lVert h(\mathbf{m^{\prime}})-h(\mathbf{m^{\prime\prime}})\right\rVert$ is an Euclidean pseudometric on $\mathcal{M}$ to $\mathbb{R}^{d}$. Since the set $\mathcal{Y}$ of all Euclidean pseudometrics from $\mathcal{M}$ to $\mathbb{R}^{d}$ is closed and complete, $\sigma((\boldsymbol{\Delta}^{(\infty)},\mathcal{P}),h)$ can be minimized by varying $h$ over the set of all Borel-measurable embedding functions from $\mathcal{M}\to\mathbb{R}^{d}$.
@@ 3 Description of the setting
A generative model is a random map from an input space or query space to an output space. Let $f_{1},\dots f_{n}$ be $n$ generative models with the shared query space $\mathcal{Q}$ and the shared output space $\mathcal{X}$, and suppose $\{q_{1},\dots q_{m}\}$ is a collection of $m$ queries. In our setting, every model $f_{i}$ responds to every query $q_{j}$ exactly $r$ times and we let $f_{i}(q_{j})_{k}$ denote the $k$-th replicate of the response of $f_{i}$ to $q_{j}$. We assume that there exists an embedding function $g$ that maps a response to a vector in $\mathbb{R}^{s}$. We let $F_{ij}$ denote the probability distribution of the vector-embedded responses of $f_{i}$ to $q_{j}$, that is, $g(f_{i}(q_{j})_{k})\sim^{iid}F_{ij}$ for all $i,j,k$.
For every $i$, the model $f_{i}$ is represented by the matrix $\bar{\mathbf{X}}_{i}\in\mathbb{R}^{m\times s}$ whose $j^{th}$ row is the mean over replicates of the $i^{th}$ model’s response to the $j^{th}$ query, that is: $(\bar{\mathbf{X}_{i}})_{j\cdot}=\frac{1}{r}\sum_{k=1}^{r}g(f_{i}(q_{j})_{k})\approx(\boldsymbol{\mu}_{i})_{j\cdot}\equiv\mathbb{E}_{F_{ij}}\left[g(f_{i}(q_{j})\right]$.
To capture the sample pairwise differences between the models, we define the matrix $\mathbf{D}\in\mathbb{R}^{n\times n}$ with entries $\mathbf{D}_{ii^{\prime}}=\frac{1}{m}||\bar{\mathbf{X}}_{i}-\bar{\mathbf{X}}_{i^{\prime}}||_{F}$. Note that, for fixed $n$ and $m$, as $r\to\infty$, $\mathbf{D}_{ii^{\prime}}$ converges to the quantity $\boldsymbol{\Delta}_{ii^{\prime}}=\frac{1}{m}||\boldsymbol{\mu}_{i}-\boldsymbol{\mu}_{i^{\prime}}||_{F}$. The matrix $\boldsymbol{\Delta}=\left(\boldsymbol{\Delta}_{ii^{\prime}}\right)_{i,i^{\prime}=1}^{n}$ is known as the model mean discrepancy matrix.
@@ 4.1 Fixed set of models and fixed set of queries
Theorem 1 states that by prompting each model with each query with enough replicates $r$, the perspective space obtained via the raw stress criterion is close to the true-but-unknown perspective space for the models with respect to $\{q_{j}\}$.
@@ 4.2 Fixed set of models and growing set of queries
For some $q\in\mathbb{N}$, there exist vectors $\boldsymbol{\phi}_{1},\dots\boldsymbol{\phi}_{n}\in\mathbb{R}^{q}$, such that for every pair $(i,i^{\prime})\in[n]^{2}$, $\boldsymbol{\Delta}_{ii^{\prime}}=\frac{1}{m}||\boldsymbol{\mu}_{i}-\boldsymbol{\mu}_{i^{\prime}}||\to\left\lVert\boldsymbol{\phi}_{i}-\boldsymbol{\phi}_{i^{\prime}}\right\rVert$ as $r,m\to\infty$.
@@
Corollary 1 states that there is a subsequence of minimizers of the raw stress criterion that is close to the true-but-unknown perspectives as the number of queries and number of replicates for each query grow. In our set up the true-but-unknown perspective space depends on the collection of models and the growing set of queries. While our results allow for arbitrary growing sets of queries, the true-but-unknown perspective space is most easily characterized when the queries come from an explicit query distribution.
@@ 4.3 Growing set of models and growing set of queries
$\int_{\mathcal{M}}\int_{\mathcal{M}}\left(\left\lVert\widehat{\boldsymbol{\psi}}_{1}^{(r_{u})}-\widehat{\boldsymbol{\psi}}_{2}^{(r_{u})}\right\rVert-\left\lVert\mathrm{mds}(\boldsymbol{\phi}_{1})-\mathrm{mds}(\boldsymbol{\phi}_{2})\right\rVert\right)^{p}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\mathcal{P}(d\boldsymbol{\phi}_{1})\mathcal{P}(d\boldsymbol{\phi}_{2})\to^{P}0\text{ as $u\to\infty$}.$
Figure 1: Numerical evidence of the consistency of $\widehat{\boldsymbol{\psi}}$ to $\boldsymbol{\psi}$ for fixed $n$ and fixed $m$ for a collection of language models (left) and a collection of text-to-image models (right). The black dots in the left figure of each pair are the 2-d perspectives of models induced with randomly selected queries, $R$ replicates each, and a domain-specific embedding function. The red circles have radius equal to the average $L_{2}$ between $\widehat{\psi}^{*}$ and model representations estimated with $r$ replicates for each query. The right figure of each pair shows the distribution of the average $L_{2}$ norm in the DKPS across models for various $r$. More replicates improves estimation quality.
@@ 5 Numerical experiments
We next provide empirical support of the consistency of the representations of models in the three settings we analyzed above. For each setting, we study a collection of large language models and a collection of text-to image models.
Language models
@@
For our language model example, we study collections of different LLaMA-3-8B-Chat [8] models. The models are parameterized by different fixed context augmentations $a_{i}$, i.e., $f_{i}=f(\;\cdot\;;a_{i})$. Each $a_{i}$ is a text string related to RA Fisher such as $a_{i}=$ “RA Fisher pioneered the principles of the design of experiments.” or $a_{i^{\prime}}=$ “RA Fisher’s view on eugenics were primarily based on anecdotes and prejudice." written by us. Given a query $q_{j}$, the base model is prompted with the appropriately formatted prompt ``$a_{i}\;q_{j}"$ $R=50$ times. In our experiments we consider up to $174$ questions about RA Fisher such as $q_{j}=$ “What is R.A. Fisher’s most well-known statistical theorem?" as queries. We consider up to $N=50$ models. The questions were sampled from ChatGPT with the prompt “Provide 200 questions related to RA Fisher". We did not include a random $26$ questions. We use the open source embedding model nomic-embed-text-v1.5 [18] to construct $\bar{\mathbf{X}}_{i}\in\mathbb{R}^{m\times 756}$ and GrasPy’s [6] implementation of multi-dimensional scaling to map the collection of $\bar{\mathbf{X}}_{i}$ to $\mathbb{R}^{2}$.
Figure 2: Numerical evidence of the consistency of $\widehat{\boldsymbol{\psi}}$ to $\boldsymbol{\psi}$ for fixed $n$ and growing $m$. The black dots in the left of each pair of figures are the 2-d perspectives of $n$ models induced with $M$ queries and $R$ replicates each. The red circles have radius equal to the average $L_{2}$ norm between the “ground truth" and model representations estimated for selected $(m,r)$ pairs. The right figures show the average maximum row $L_{2}$ norm for various $(m,r)$ pairs. More replicates and more queries improves estimation quality. The number of queries appears to have a larger effect.
Text-to-image models
For our text-to-image example, we study collections of StableDiffusion-3.5-medium [10]. As with our language model example, we parameterize different models by different fixed context augmentations. Each fixed augmentation $a_{i}$ is an instruction to produce an image in the style of a famous artist – for example, $a_{i}$ = “in the style of Pablo Picasso" or $a_{i^{\prime}}$ = “in the style of Leonardo da Vinci" – and $f_{i}=f(\;\cdot\;;a_{i})$. We consider up to $N=100$ models. Given a query $q_{j}$, the base model is prompted with the appropriately formatted prompt ``$q_{j}\;a_{i}"$ $R=25$ times. Each query $q_{j}$ is an instruction to produce an image of a noun. For example, one of the prompts for the model corresponding to da Vinci is “An apple in the style of Leonardo da Vinci". The nouns were generated from ChatGPT with the prompt “Provide 500 nouns." We use the open source embedding model nomic-embed-vision-v1.5 to map the generated images to a vector space and construct the $\bar{\mathbf{X}}_{i}\in\mathbb{R}^{m\times 768}$. We again use GrasPy’s implementation of multi-dimensional scaling to induce the data kernel perspective space of the models. To avoid sounding repetitive, we use $N$, $M$, and $R$ to mean the maximum number of models, queries, and replicates for both the language model experiments and the text-to-image experiments. For the language model experiment, $N=50$, $M=174$, and $R=50$. For the text-to-image experiment, $N=100$, $M=500$, and $R=25$.
@@
For both sets of experiments we do not have the “ground truth" DKPS $\boldsymbol{\psi}$ in any parameter setting, and thus treat $\widehat{\boldsymbol{\psi}}$ estimated with $r=R$ as a proxy for the ground truth. For settings with growing $n$ and/or growing $m$, we similarly treat $\widehat{\boldsymbol{\psi}}^{(R)}$ estimated with $n=N$ and $m=M$ as the ground truth. We refer to these estimates as $\widehat{\boldsymbol{\psi}}^{*}$. We let $d=2$ in all settings for visualization purposes. Finally, in settings where $n,m,$ or $r$ grow, we sample from the appropriate set of models, queries, or replicates with replacement to calculate $\widehat{\boldsymbol{\psi}}$ and report the average $L_{2}$ norm or average two-to-infinity norm of the difference between $\boldsymbol{\widehat{\psi}}^{*}$ and $\boldsymbol{\widehat{\psi}}\mathbf{W}$ of 10 bootstrap samples, where $\mathbf{W}\in\mathcal{O}(2)$ is the Procrustes solution [12].
@@ 5.1 Fixed collection of models and fixed set of queries
As with the theoretical analysis, our empirical analysis starts in the simplest setting where $n$ and $m$ are fixed while $r$ grows. The collection of models and set of queries were selected at random without replacement – $n=10$ and $m=20$ for the language model example, $n=20$ and $m=100$ for the text-to-image example. The target DKPS $\boldsymbol{\widehat{\psi}}^{*}$ for the two experiments are shown on the left for each pair of figures of Figure 1. Each black dot represents a model. The red circles have radius equal to the average (across bootstrap samples) $L_{2}$ norm between $\boldsymbol{\widehat{\psi}}^{*}$ and $\boldsymbol{\widehat{\psi}}^{(r)}\mathbf{W}$ for selected $r$. The right figure for each pair of figures of Figure 1 shows the distribution of the average (across bootstrap samples) $L_{2}$ norm for each model for more values of $r$. The decreasing average $L_{2}$ norm for all models as $r$ increases supports Theorem 1.